# A Novel Patch-Based TDA Approach for Computed Tomography

### v1

@@ 1 Introduction
The development of computer aided diagnostic tools has gained significant interest in recent years, particularly in the prognostication of cancer where they offer the potential to derive non-invasive biomarkers from various medical imaging technologies. Computed tomography (CT) scans are, in particular, a primary source of data to build robust machine learning (ML) models since they provide rich insights about the disease through their volumetric nature. Leveraging this comprehensive imaging modality, many ML algorithms have been proposed across the last decade, and new ones continue to appear with alarming frequency. Classical ML approaches such as radiomics or sophisticated deep learning-based approaches have shown success in numerous tasks. However, radiomic features, which rely on pixel-wise comparisons, are often sensitive to variations in image acquisition settings, for instance differences in resolution or contrast [mayerhoefer2020introduction, lee2021radiomics, rizzo2018radiomics], and deep learning methods, while addressing some of these challenges, are criticized for insufficient explainability, due to opaque decision-making, and their high computational demands hinder their accessibility [hayashi2019right, vial2018role, majeed2020issues]. Topological data analysis (TDA) has emerged as a promising approach, to address some of these limitations, providing a unique way to extract richer patterns and insight from data. TDA is built mainly upon theoretical concepts from a field of mathematics called algebraic topology, analyzing the topological structure of the data across different scales, revealing complex relationships and uncovering hidden patterns that may not be visible to the naked eye [carlsson2009topology, lum2013extracting].
@@
When it comes to developing ML models on medical imaging data with volumetric structure (e.g., 3D CT scans), employing robust feature engineering approaches is vital, particularly one that take advantage of the temporal information presents in the data. A popular choice in the realm of TDA to extract topological features by constructing persistent homology (PH) from such data is 3D cubical complex filtration [somasundaram2021persistent, tanabe2021homological]. Although this approach has shown success on grid structured data, such as 2D images, it may not provide the optimal result and be time-efficient on high resolution volumetric data. In this study, we propose a patch-based PH construction method to address these challenges. This approach mainly concentrates on the transformation of the volumetric data, with specific focus on 3D CT scans modalities in this study, to an $d$-dimensional point cloud through the process of 3D image patch summarization to $d$-dimensional point. With a point cloud representing the 3D image, we have utilized alpha complex filtration, an efficient approach to construct PH from point cloud data. Note that another advantage of this approach is that, unlike cubical complex, with point cloud data higher dimensional structures beyond connected components, loops and enclosed voids can be explored. In our previous work [ali2026a], a PCA-based approach was proposed which only utilized PCA for patch intensities summarization. However, this study provides a more comprehensive technique that employs different patch intensities summarization approaches and also integrates compressed patch coordinates representation into the pipeline outperforming the previous approach.
@@
TDA has demonstrated significant success in the development of models based on different medical imaging modalities such as X-ray, CT, MRI, and ultrasound [singh2023topological]. For instance, in the assessment of tumour heterogeneity in lung adenocarcinomas based on CT images [kawata2021representation], analysis of osteoarthritis from MRI images [pedoia2018mri], classification of chronic obstructive pulmonary disease and explaining the geometric structure of the airway system through lung CT images [belchi2018lung], evaluation of the accuracy of hepatic cancer prediction via MRI images [oyama2019hepatic], distinguishing squamous cell carcinoma from adenocarcinoma and benign from malignant tumours in lung CT images [vandaele2023topological], unveiling the dynamical organization of the brain using fMRI data [saggar2018towards] and characterization of the 3D structure of trabecular bone and predicting bone strength using bone MR images [gomberg2000topological]. To the best of our knowledge, the proposed technique is one of the first alternative approaches to cubical complex-based PH computation for volumetric medical imaging data.
@@ 2.1 Topological Data Analysis
Topology is a field of mathematics concerned with the connectivity and continuity properties of objects. An important aspect of the topological spaces and their properties is their ability to be coordinate free, invariant under continuous deformation and to be expressed in compressed form [lum2013extracting]. Using tools in the field of TDA we can extract topological invariants of complex and high dimensional data which are stable in terms of small perturbations to input noise and can be employed with various ML tasks such as classification. Topological invariant examples studied in this work are connected components, loops and cavities (i.e. enclosed voids).
@@
These topological invariants can be quantified by what are called Betti numbers. The first $N$ Betti numbers are used to differentiate objects with differing topology in $N$-dimensional space. For instance, the Betti number of 2D objects includes $(\beta_{0},\beta_{1})$, and $(\beta_{0},\beta_{1},\beta_{2})$ for 3D, where $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ refer to connected components, loops and voids respectively. Although this characterization is simple to describe in abstraction, its practical computation requires a smart machinery. This can be achieved through homology theory, a well-established technique in algebraic topology. To apply homology theory to compute Betti numbers of objects, they must be represented in combinatorial structures such as simplicial or cubical complexes [otter2017roadmap].
@@
PH is a principal tool in the field of TDA, used to quantify the shape of data at different scales (computing Betti numbers), tracking the evolution of topological features such as connected components, cycles and cavities. Homology provides the theoretical framework for persistent homology. In order to compute topological features from data, one need to utilize a combinatorial approach, known as filtration, such as a simplicial or cubical complex, to build PH. A simplicial complex (SC) is a topological space build from a set of multi-dimensional triangular subsets of $\mathbb{R}^{n}$, by joining some of them together along their boundary components such as nodes, edges and faces. SCs transform the continuous domain of topological spaces to the discrete domain of data [nanda2013simplicial]. Figure 1 demonstrate this process on a sample point cloud data. The evolution of topological structures, including connected components and loops can be observed from the figure as the filtration process advances from left to right, by increasing the distance threshold. Through this process, each topological feature is tracked using a horizontal bar, where the start of the bar indicates when the feature first appears, and the end of the bar indicates when the feature disappears or merges with another feature. A collection of such bars is called a persistence barcode (PB). The concept of PBs will be explained further in later sections. In general, the process of acquiring topological features from data involves two main steps. First constructing PH from the data through a filtration method, to obtain PBs and second, featurizing the space of PBs using vectorization methods to be ready for ML algorithms or for further analysis. In this work, we utilize two distinct filtration approaches: alpha complex, for building PH from point cloud data, and the cubical complex which is tailored for grid-structured data such as digital images. In the following sections, these methods will be explained in detail.
@@ 2.2 Alpha Complex
Let $X=\bigcup_{i\in I}U_{i}$ be a topological space. Then, the cover of $X$ is defined as a non-empty and finite collection of open sets $\mathcal{U}=\{U_{i}\}_{i\in I}$. More specifically, $\mathcal{U}$ is called a good cover if all individual sets $U_{i}\in\mathcal{U}$ and all finite, non-empty intersections of those sets are contractible, i.e., they can be continuously deformed into a single point [carlsson2014topological, bott2013differential].
@@
The alpha complex $\mathcal{A}_{\varepsilon}(X)$ is the nerve of the covers shaped by $\mathbb{R}_{\varepsilon}(x)$ for all $x\in X$, [pun2022persistent] i.e.
@@ 2.3 Cubical Complex
Cubical complex is a combinatorial approach to construct PH from grid-structured data, for example, 2D or 3D digital images. Cubical complex is constructed by ordering the cells (vertices, edges, squares, etc.) using a scalar function, voxel intensity value in the context of 3D CT scan images where cells are added to the complex based on the ascending order of their voxel intensity values. Mathematically, a finite cubical complex in $\mathbb{R}^{3}$ is a collection of cubes aligned on the grid $G^{3}$, satisfying some conditions analogous to a SC [garin2019topological]. A $3$-dimensional image can be represented as a map $\eta:I\subseteq G^{3}\rightarrow\mathbb{R}$. For an element $v\in I$, called a voxel (pixel when $d=2$), its intensity or grayscale value is given by $\eta(v)$. There are several ways to represent digital images as cubical complexes. A common construction is to represent each voxel as a top-dimensional $3$-cube and include all of its faces in the complex. A 3D CT scan comes with a natural filtration within the grayscale values of its voxels. A function is obtained on the resulting cubical complex $C$ as follows by extending the values of voxels to all the cubes $\omega\in C$:
@@
be the $i$-th sublevel set of $C$. This yields a filtration of the cubical complex defined as $\{C_{i}\}_{i\in Im(I)}$ and indexed by the values of the function $\eta$. [garin2019topological].
@@ 2.4 Persistence Barcodes
where $B(p,q)$ corresponds to a bar with birth at $p$ and death at $q$. Note that a barcode has a multiset structure, meaning that an element $(p,q)$ in the barcode $B$ may occur multiple times [hofer2019learning]. Therefore, the barcode $B$ has a multiplicity function $\mu:B\rightarrow\mathbb{Z}_{>0}$.
@@ 2.5 Vectorization
The PBs acquired via a filtration process (here alpha or cubical complex), are not directly compatible with ML algorithms since they are not easily amenable to statistical analysis; for instance, their Fréchet mean is not unique [hofer2019learning]. To overcome this bottleneck, barcodes are often embedded into vector spaces (e.g., Hilbert or Banach spaces), where well-defined inner products enable standard ML operations. Several vectorization techniques have been proposed in recent years to transform PBs into suitable feature vectors [chung2022persistence, chevyrev2018persistence, berry2020functional, chintakunta2015entropy]. Persistent statistical vectorization is a frequently used approach to featurize the space of PBs which provides a compact summary of PBs and has shown robustness and success in comparison with other vectorization approaches to analysis digital images [ali2023survey]. This approach constructs a feature vector by computing different statistical quantities from a given PB including the mean, median, standard deviation, full range, inter-quartile range, and the 10th, 25th, 75th, and 90th percentiles of the birth $p$, the death $q$, the lifespans qpq-p and the midpoints of the bar intervals $[p,q]$ in a PB. Entropy and bar counts are also included to further extend the feature set. The entropy of $\mu$, introduced in [chintakunta2015entropy], is a real number defined as follows:
@@ 2.6 Patch-based TDA: The Proposed Approach
For the coordinate encoding part, after preliminary experiments with a number of methods such as Morton code ($Z$-order curve) and $L_{2}$ norm, the Morton code algorithms [morton1966computer] is utilized to compress the three coordinates ($x$, $y$ and $z$), corresponding to the approximate centre of the patch, to a single value. In parallel, for the second part of the pipeline, intensities encoding, two distinct approaches are considered: 1) employing PCA to reduce the dimension of the flattened patch and consequently utilizing $d$ PCA components to create the point cloud, 2) computing statistical quantities (stats), for instance mean, mode and median, from the flattened patch to create the final point where each one of these stats corresponds to an axis in the space. This procedure is demonstrated in Figure 4, part A. Finally, applying the patch-to-point transformation techniques on all the patches taken from a volumetric ROI of a 3D image, a point cloud is obtained as depicted in Figure 4, part B. It is worth mentioning that, during this process only cubic patches with at least one voxel corresponding to the ROI are considered and converted to points contributing to the final point cloud, i.e, void or empty patches are discarded. This step is crucial to enhance the downstream computation of PH from the point cloud.
@@
There are different ways to compute PH from a point cloud. In this study, alpha complex filtration [edelsbrunner1983shape, boissonnat2018geometric] is utilized to compute PH in dimensions zero, one and two, corresponding to the number of connected components, loops, and cavities, respectively. Hence, obtaining three persistence barcodes. This follows by the vectorization process, computing persistent statistical features for each barcode, yielding three feature vectors. Finally, these feature vectors are concatenated and used as input to a ML classifier. This workflow is illustrated in Figure 5.
@@ 2.7 Datasets
In this work, four diverse datasets of CT scans are utilized in order to conduct a comprehensive assessment of the patch-based TDA approach and benchmark its performance. These include the publicly available dataset of kidney tumours KiTS19 [heller2019kits19], the FLARE22 dataset from MICCAI FLARE22 challenge [Ma2023-vo, ma2021abdomenct] which is a collection of abdominal CT images, and two in-house datasets. Sample CT slices of different classes of each dataset is illustrated in Figure 6. Table 1 presents a summary of key details about each dataset.
@@ 3 Results
Pancreas
Tumours
Stats $6\!\times\!6\!\times\!6$ iqr, entropy LR 75.16 ± 8.97 73.05 ± 9.25 60 ± 20 81.15 ± 10.57 57.22 ± 14.94
@@
Pancreas
Tumours
Cubical Complex LR 55.03 ± 11.84 58.44 ± 9.85 68 ± 17.89 49.87 ± 18.91 45.9 ± 8.54
@@
KiTS19
$57\!\times\!220\!\times\!220$ Patch-Based $6\!\times\!6\!\times\!6$ mean, median, range 4D 0.26
@@
Pancreas
Tumours
$33\!\times\!120\!\times\!120$ Patch-Based $6\!\times\!6\!\times\!6$ iqr, entropy 3D 0.07
@@
CRLM
$87\!\times\!323\!\times\!400$ Patch-Based $5\!\times\!5\!\times\!5$ median, entropy, max 4D 31.69

### v2

@@ 1 Introduction
The development of computer aided diagnostic tools has gained significant interest in recent years, particularly in the prognostication of cancer where they offer the potential to derive non-invasive biomarkers from various medical imaging technologies. Computed tomography (CT) scans are, in particular, a primary source of data to build robust machine learning (ML) models since they provide rich insights about the disease through their volumetric nature. Leveraging this comprehensive imaging modality, many ML algorithms have been proposed across the last decade, and new ones continue to appear with alarming frequency. Classical ML approaches such as radiomics or sophisticated deep learning-based approaches have shown success in numerous tasks. However, radiomic features, which rely on pixel-wise comparisons, are often sensitive to variations in image acquisition settings, for instance differences in resolution or contrast [25, 20, 32], and deep learning methods, while addressing some of these challenges, are criticized for insufficient explainability, due to opaque decision-making, and their high computational demands hinder their accessibility [16, 38, 24]. Topological data analysis (TDA) has emerged as a promising approach, to address some of these limitations, providing a unique way to extract richer patterns and insight from data. TDA is built mainly upon theoretical concepts from a field of mathematics called algebraic topology, analyzing the topological structure of the data across different scales, revealing complex relationships and uncovering hidden patterns that may not be visible to the naked eye [7, 21].
@@
When it comes to developing ML models on medical imaging data with volumetric structure (e.g., 3D CT scans), employing robust feature engineering approaches is vital, particularly one that take advantage of the temporal information presents in the data. A popular choice in the realm of TDA to extract topological features by constructing persistent homology (PH) from such data is 3D cubical complex filtration [35, 36]. Although this approach has shown success on grid structured data, such as 2D images, it may not provide the optimal result and be time-efficient on high resolution volumetric data. In this study, we propose a patch-based PH construction method to address these challenges. This approach mainly concentrates on the transformation of the volumetric data, with specific focus on 3D CT scans modalities in this study, to an $d$-dimensional point cloud through the process of 3D image patch summarization to $d$-dimensional point. With a point cloud representing the 3D image, we have utilized alpha complex filtration, an efficient approach to construct PH from point cloud data. Note that another advantage of this approach is that, unlike cubical complex, with point cloud data higher dimensional structures beyond connected components, loops and enclosed voids can be explored. In our previous work [1], a PCA-based approach was proposed which only utilized PCA for patch intensities summarization. However, this study provides a more comprehensive technique that employs different patch intensities summarization approaches and also integrates compressed patch coordinates representation into the pipeline outperforming the previous approach.
@@
TDA has demonstrated significant success in the development of models based on different medical imaging modalities such as X-ray, CT, MRI, and ultrasound [34]. For instance, in the assessment of tumour heterogeneity in lung adenocarcinomas based on CT images [19], analysis of osteoarthritis from MRI images [30], classification of chronic obstructive pulmonary disease and explaining the geometric structure of the airway system through lung CT images [3], evaluation of the accuracy of hepatic cancer prediction via MRI images [29], distinguishing squamous cell carcinoma from adenocarcinoma and benign from malignant tumours in lung CT images [37], unveiling the dynamical organization of the brain using fMRI data [33] and characterization of the 3D structure of trabecular bone and predicting bone strength using bone MR images [14]. To the best of our knowledge, the proposed technique is one of the first alternative approaches to cubical complex-based PH computation for volumetric medical imaging data.
@@ 2.1 Topological Data Analysis
Topology is a field of mathematics concerned with the connectivity and continuity properties of objects. An important aspect of the topological spaces and their properties is their ability to be coordinate free, invariant under continuous deformation and to be expressed in compressed form [21]. Using tools in the field of TDA we can extract topological invariants of complex and high dimensional data which are stable in terms of small perturbations to input noise and can be employed with various ML tasks such as classification. Topological invariant examples studied in this work are connected components, loops and cavities (i.e. enclosed voids).
@@
These topological invariants can be quantified by what are called Betti numbers. The first $N$ Betti numbers are used to differentiate objects with differing topology in $N$-dimensional space. For instance, the Betti number of 2D objects includes $(\beta_{0},\beta_{1})$, and $(\beta_{0},\beta_{1},\beta_{2})$ for 3D, where $\beta_{0}$, $\beta_{1}$ and $\beta_{2}$ refer to connected components, loops and voids respectively. Although this characterization is simple to describe in abstraction, its practical computation requires a smart machinery. This can be achieved through homology theory, a well-established technique in algebraic topology. To apply homology theory to compute Betti numbers of objects, they must be represented in combinatorial structures such as simplicial or cubical complexes [28].
@@
PH is a principal tool in the field of TDA, used to quantify the shape of data at different scales (computing Betti numbers), tracking the evolution of topological features such as connected components, cycles and cavities. Homology provides the theoretical framework for persistent homology. In order to compute topological features from data, one need to utilize a combinatorial approach, known as filtration, such as a simplicial or cubical complex, to build PH. A simplicial complex (SC) is a topological space build from a set of multi-dimensional triangular subsets of $\mathbb{R}^{n}$, by joining some of them together along their boundary components such as nodes, edges and faces. SCs transform the continuous domain of topological spaces to the discrete domain of data [27]. Figure 1 demonstrate this process on a sample point cloud data. The evolution of topological structures, including connected components and loops can be observed from the figure as the filtration process advances from left to right, by increasing the distance threshold. Through this process, each topological feature is tracked using a horizontal bar, where the start of the bar indicates when the feature first appears, and the end of the bar indicates when the feature disappears or merges with another feature. A collection of such bars is called a persistence barcode (PB). The concept of PBs will be explained further in later sections. In general, the process of acquiring topological features from data involves two main steps. First constructing PH from the data through a filtration method, to obtain PBs and second, featurizing the space of PBs using vectorization methods to be ready for ML algorithms or for further analysis. In this work, we utilize two distinct filtration approaches: alpha complex, for building PH from point cloud data, and the cubical complex which is tailored for grid-structured data such as digital images. In the following sections, these methods will be explained in detail.
@@ 2.2 Alpha Complex
Let $X=\bigcup_{i\in I}U_{i}$ be a topological space. Then, the cover of $X$ is defined as a non-empty and finite collection of open sets $\mathcal{U}=\{U_{i}\}_{i\in I}$. More specifically, $\mathcal{U}$ is called a good cover if all individual sets $U_{i}\in\mathcal{U}$ and all finite, non-empty intersections of those sets are contractible, i.e., they can be continuously deformed into a single point [8, 6].
@@
The alpha complex $\mathcal{A}_{\varepsilon}(X)$ is the nerve of the covers shaped by $\mathbb{R}_{\varepsilon}(x)$ for all $x\in X$, [31] i.e.
@@ 2.3 Cubical Complex
Cubical complex is a combinatorial approach to construct PH from grid-structured data, for example, 2D or 3D digital images. Cubical complex is constructed by ordering the cells (vertices, edges, squares, etc.) using a scalar function, voxel intensity value in the context of 3D CT scan images where cells are added to the complex based on the ascending order of their voxel intensity values. Mathematically, a finite cubical complex in $\mathbb{R}^{3}$ is a collection of cubes aligned on the grid $G^{3}$, satisfying some conditions analogous to a SC [13]. A $3$-dimensional image can be represented as a map $\eta:I\subseteq G^{3}\rightarrow\mathbb{R}$. For an element $v\in I$, called a voxel (pixel when $d=2$), its intensity or grayscale value is given by $\eta(v)$. There are several ways to represent digital images as cubical complexes. A common construction is to represent each voxel as a top-dimensional $3$-cube and include all of its faces in the complex. A 3D CT scan comes with a natural filtration within the grayscale values of its voxels. A function is obtained on the resulting cubical complex $C$ as follows by extending the values of voxels to all the cubes $\omega\in C$:
@@
be the $i$-th sublevel set of $C$. This yields a filtration of the cubical complex defined as $\{C_{i}\}_{i\in Im(I)}$ and indexed by the values of the function $\eta$. [13].
@@ 2.4 Persistence Barcodes
where $B(p,q)$ corresponds to a bar with birth at $p$ and death at $q$. Note that a barcode has a multiset structure, meaning that an element $(p,q)$ in the barcode $B$ may occur multiple times [18]. Therefore, the barcode $B$ has a multiplicity function $\mu:B\rightarrow\mathbb{Z}_{>0}$.
@@ 2.5 Vectorization
The PBs acquired via a filtration process (here alpha or cubical complex), are not directly compatible with ML algorithms since they are not easily amenable to statistical analysis; for instance, their Fréchet mean is not unique [18]. To overcome this bottleneck, barcodes are often embedded into vector spaces (e.g., Hilbert or Banach spaces), where well-defined inner products enable standard ML operations. Several vectorization techniques have been proposed in recent years to transform PBs into suitable feature vectors [11, 9, 4, 10]. Persistent statistical vectorization is a frequently used approach to featurize the space of PBs which provides a compact summary of PBs and has shown robustness and success in comparison with other vectorization approaches to analysis digital images [2]. This approach constructs a feature vector by computing different statistical quantities from a given PB including the mean, median, standard deviation, full range, inter-quartile range, and the 10th, 25th, 75th, and 90th percentiles of the birth $p$, the death $q$, the lifespans qpq-p and the midpoints of the bar intervals $[p,q]$ in a PB. Entropy and bar counts are also included to further extend the feature set. The entropy of $\mu$, introduced in [10], is a real number defined as follows:
@@ 2.6 Patch-based TDA: The Proposed Approach
For the coordinate encoding part, after preliminary experiments with a number of methods such as Morton code ($Z$-order curve) and $L_{2}$ norm, the Morton code algorithms [26] is utilized to compress the three coordinates ($x$, $y$ and $z$), corresponding to the approximate centre of the patch, to a single value. In parallel, for the second part of the pipeline, intensities encoding, two distinct approaches are considered: 1) employing PCA to reduce the dimension of the flattened patch and consequently utilizing $d$ PCA components to create the point cloud, 2) computing statistical quantities (stats), for instance mean, mode and median, from the flattened patch to create the final point where each one of these stats corresponds to an axis in the space. This procedure is demonstrated in Figure 4, part A. Finally, applying the patch-to-point transformation techniques on all the patches taken from a volumetric ROI of a 3D image, a point cloud is obtained as depicted in Figure 4, part B. It is worth mentioning that, during this process only cubic patches with at least one voxel corresponding to the ROI are considered and converted to points contributing to the final point cloud, i.e, void or empty patches are discarded. This step is crucial to enhance the downstream computation of PH from the point cloud.
@@
There are different ways to compute PH from a point cloud. In this study, alpha complex filtration [12, 5] is utilized to compute PH in dimensions zero, one and two, corresponding to the number of connected components, loops, and cavities, respectively. Hence, obtaining three persistence barcodes. This follows by the vectorization process, computing persistent statistical features for each barcode, yielding three feature vectors. Finally, these feature vectors are concatenated and used as input to a ML classifier. This workflow is illustrated in Figure 5.
@@ 2.7 Datasets
In this work, four diverse datasets of CT scans are utilized in order to conduct a comprehensive assessment of the patch-based TDA approach and benchmark its performance. These include the publicly available dataset of kidney tumours KiTS19 [17], the FLARE22 dataset from MICCAI FLARE22 challenge [23, 22] which is a collection of abdominal CT images, and two in-house datasets. Sample CT slices of different classes of each dataset is illustrated in Figure 6. Table 1 presents a summary of key details about each dataset.
@@ 3 Results
Pancreas Tumours Stats $6\!\times\!6\!\times\!6$ iqr, entropy LR 75.16 ± 8.97 73.05 ± 9.25 60 ± 20 81.15 ± 10.57 57.22 ± 14.94
@@
Pancreas Tumours Cubical Complex LR 55.03 ± 11.84 58.44 ± 9.85 68 ± 17.89 49.87 ± 18.91 45.9 ± 8.54
@@
KiTS19 $57\!\times\!220\!\times\!220$ Patch-Based $6\!\times\!6\!\times\!6$ mean, median, range 4D 0.26
@@
Pancreas Tumours $33\!\times\!120\!\times\!120$ Patch-Based $6\!\times\!6\!\times\!6$ iqr, entropy 3D 0.07
@@
CRLM $87\!\times\!323\!\times\!400$ Patch-Based $5\!\times\!5\!\times\!5$ median, entropy, max 4D 31.69